OpenStudy (fredfredburgeryes):

The step by step solve problem Help me understand step by step how to divide this following monstrosity? will fan and medal! @phi (8x^4y^3 + 4x^3y^2 - 2x^2y - 12x^2y^4) divided by -2x^2y

OpenStudy (phi):

first, did you know when you subtract fractions like \[ \frac{10}{20} - \frac{5}{20} \] you have the same denominator so you do this \[ \frac{10}{20} - \frac{5}{20}= \frac{10-5}{20} \] ?

OpenStudy (fredfredburgeryes):

yes, but is this not some sort of polynomial division or something?

OpenStudy (phi):

poly means many (in Greek I think) you have only one term, so it's not very poly.. (its a monomial ... mono means 1) though you could call it a polynomial... most people would not bother)

OpenStudy (fredfredburgeryes):

I am just curious as to the step by step on how to solve this thing xD I learn best from visual references and text

OpenStudy (phi):

the point is that the fraction idea works the other way: \[ \frac{10-5}{20}=\frac{10}{20} - \frac{5}{20} \] and that idea works for your problem. in other words, notice your denominator you have a -2 that means (if it divides evenly) you should divide -2 into each term up top. to get a "new top" can you do that?

OpenStudy (fredfredburgeryes):

so divide each term above by -2? like 8x^4y^3/-2 and so on?

OpenStudy (phi):

yes. we are doing this in small steps

OpenStudy (phi):

what did you get?

OpenStudy (fredfredburgeryes):

-4x^4y^3?

OpenStudy (phi):

notice that \[ \frac{8 x^4 y^3}{-2}= \frac{8}{-2} x^4 y^3 \]

OpenStudy (phi):

yes 8 divided by -2 is -4 so you get -4 * x^4 * y^3 for the first term. but you also have to divide -2 into all the terms in the top.

OpenStudy (fredfredburgeryes):

so 4x^3y^2/-2 would be -2x^3y^2?

OpenStudy (fredfredburgeryes):

and we are dividing each by not -2 but the whole denominator?

OpenStudy (phi):

after you get the new top you now have \[ \frac{-4x^4y^3 -2x^3y^2+x^2y + 6x^2y^4}{x^2y}\]

OpenStudy (phi):

now you can divide each term by y remember that , for example y^3 means y*y*y so when you divide y*y*y by y you get \[ \frac{y}{y} \cdot y \cdot y\] which is 1*y*y or \(y^2\) most people do it the"fast way" and do "top exponent" minus bottom exponent in other words y^3 / y^1 simplifies to y^2

OpenStudy (fredfredburgeryes):

so then that changes -4x^4y^3 into -4x^4y^2?

OpenStudy (phi):

yes. now do the other 3 terms that are up top

OpenStudy (fredfredburgeryes):

y^2/y^1 should be y correct? and then y^4/y^1=y^3?

OpenStudy (fredfredburgeryes):

@phi what next?

OpenStudy (phi):

ok. what do you have for the top now ?

OpenStudy (fredfredburgeryes):

-4x^4y^2 - 2x^3y - x^2y + 6x^2y^3

OpenStudy (phi):

the third term is not correct. first, it should be positive, because you did -2/-2 and that is +1 then you should have done y/y and that is 1

OpenStudy (fredfredburgeryes):

so it would be cancelled out?

OpenStudy (phi):

\[ \frac{-4x^4y^2 -2x^3y^1+x^2y^0 + 6x^2y^3}{x^2} \] I make the y's exponent in the third term y^(1-1) = y^0 anything to the zero power is 1 (which is another way to do y/y = 1). anyway, yes it "cancels",

OpenStudy (phi):

now divide x*x into each term up top. the fast way: top exponent minus 2 (2 is the bottom x's exponent) the simple way: if you have x*x*x*x and you divide by x*x two of the x's are cancelled. either way, you get x*x or x^2 up top in the first term

OpenStudy (fredfredburgeryes):

-4x^2y^2 - 2x3y + x2y + 6xy^3?

OpenStudy (phi):

what did you do? so far we have \[ \frac{-4x^4y^2 -2x^3y^1+x^2 + 6x^2y^3}{x^2} \] remember the x^2 is divided into each term up top

OpenStudy (fredfredburgeryes):

o lol 1sec

OpenStudy (fredfredburgeryes):

confused now xD the last term didn't come up right so far I have -4x^2y^2 - 2xy + y + 6xy^3

OpenStudy (phi):

the 3rd term was originally - 2x^2y and you divide that by the original -2x^2 y notice you do *not* get y

OpenStudy (phi):

the fourth (last) term is \[ \frac{6x^2y^3}{x^2} \] if you have x*x up top divided by x*x below, you get x/x * x/x or 1*1 or 1 i.e. the x^2 is "cancelled" by the x^2 down below

OpenStudy (fredfredburgeryes):

so itd be 6y^3?

OpenStudy (phi):

for the 4th term, yes what do you get for the third term?

OpenStudy (fredfredburgeryes):

didn't we make the 3rd term a positive tho?

OpenStudy (phi):

yes, it will be positive.

OpenStudy (fredfredburgeryes):

so I have to solve x^2y/x^2?

OpenStudy (phi):

not exactly. re-do that one from scratch: \[ \frac{-2 x^2 y}{-2 x^2 y} \]

OpenStudy (fredfredburgeryes):

-1?

OpenStudy (phi):

-2/-2 is not -1

OpenStudy (phi):

minus divided by minus is plus

OpenStudy (fredfredburgeryes):

1

OpenStudy (phi):

yes. so the answer is \[ -4x^2y^2 -2xy+1 + 6y^3 \]

OpenStudy (phi):

though I might write it in this order \[ 6y^3 -4x^2y^2 -2xy+1\]

OpenStudy (fredfredburgeryes):

what kind of polynomial is it?

OpenStudy (phi):

find the "highest degree" term by adding up the exponents of the variables in each term

OpenStudy (fredfredburgeryes):

4th degree polynomial?

OpenStudy (phi):

for example, 6y^3 has just 3 as the exponent -4x^2 y^2 has 2+2 = 4 -2xy has 1+1 (x^1 * y^1) so 2 and +1 has no variables so 0 the highest degree term is 4 so it's a 4th degree polynomial

OpenStudy (fredfredburgeryes):

Thanks!